OpenStudy (anonymous):

what the integration of 1/(1-x^2)*(4+tanh"inverse"x)^(1/2)

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\frac{1}{1-x^2}(4+\tanh^{-1}x)~dx}\) like this?

OpenStudy (solomonzelman):

use u substitution, \(\large\color{slate}{\displaystyle u=4+\tanh^{-1}x}\)

OpenStudy (misty1212):

i bet it is in the denominator

OpenStudy (anonymous):

no its integaration 1/((1-x^2)*(4+tanh"inverse"x)^(1/2))

OpenStudy (misty1212):

that way you can make \(u=arctanh(x)\) and do a u-sub in one step i could be wrong

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\frac{1}{(1-x^2)(4+\tanh^{-1}x)}~dx}\) still do the same u substitution

OpenStudy (anonymous):

yes give me the full answer to make sure that i correctly solve

OpenStudy (solomonzelman):

we don't give answers, or at least try not to. We can help you out though...

OpenStudy (misty1212):

\[\int\frac{dx}{(1-x^2)\sqrt{(4-tanh^{-1}(x)}}\]

OpenStudy (misty1212):

just a guess, maybe it is something else

OpenStudy (solomonzelman):

Your problem is? \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\frac{1}{(1-x^2)(4+\tanh^{-1}x)}~dx}\) But no matter where \(\large\color{slate}{\displaystyle 4+\tanh^{-1}x}\) in the root, just in denominator, or multiplied times the fraction, Do: \(\large\color{blue}{\displaystyle u=4+\tanh^{-1}x}\)

OpenStudy (anonymous):

thank you all for helping

OpenStudy (solomonzelman):

(I am still not clear about what is exactly the problem )

OpenStudy (solomonzelman):

is it the last thing I suggested?

OpenStudy (misty1212):

one more guess \[\huge \int\frac{dx}{(1-x^2)\sqrt{(4+tanh^{-1}(x)}}\]

OpenStudy (anonymous):

i solved the problem thanks all

OpenStudy (solomonzelman):

what was the problem? what was the answer you got?

OpenStudy (anonymous):